OpenStudy (anonymous):

find h' (x), where f(x) is an unspecified differentiable function. Use the notation f' to denote the derivative of f Example : if h(x)=4(x)^2 then h'(x)=4(2f(x)f'(x))=8f(x)f'(x) a) h(x)=3f(x)^4 B) h(x)=6x^8f(x) C)h(x)=(1-e^x)^3/f(x) D) H(x)=5sqrtf(x)^4

zepdrix (zepdrix):

\[\Large\bf\sf h(x)\quad=\quad 3f(x)^4\]Mmmm so what are you having trouble with?

zepdrix (zepdrix):

Function notation can be hard to get comfortable with. f(x) just means that the right side contains a bunch of `stuff` involving x. They didn't specify exactly what type of stuff. But we can apply the rules that we know.

zepdrix (zepdrix):

\[\Large\bf\sf h(x)\quad=\quad 3(stuff)^4\]\[\Large\bf\sf h'(x)\quad=\quad 3\cdot 4(stuff)^3(stuff)'\]Power rule, then chain rule, yes?

zepdrix (zepdrix):

\[\Large\bf\sf h(x)\quad=\quad 3f(x)^4\]\[\Large\bf\sf h'(x)\quad=\quad 3\cdot4f(x)^3\cdot f'(x)\]

zepdrix (zepdrix):

Where you at Louis!! :L

OpenStudy (anonymous):

confused:(

OpenStudy (anonymous):

do i plug in h(x)=3f(x)^4 into H'

zepdrix (zepdrix):

The second line that I wrote, that derivative, that one is done. Part a) is done. You don't go any further than that. (You can multiply the 3 and 4 together I guess).

zepdrix (zepdrix):

Are you confused how I got that line?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

so it will be \[12f(x)f'(x)\]

zepdrix (zepdrix):

Yes.

OpenStudy (anonymous):

ok i was reading it wrong

zepdrix (zepdrix):

Ah :3

zepdrix (zepdrix):

Understand how to do the next one? Looks like we have the `product` of two x-things.

OpenStudy (anonymous):

no not really but i'm trying to work it out now

zepdrix (zepdrix):

If you have a `product`, you use a certain `ruuuuuuule` :X

OpenStudy (anonymous):

will it start of 8x^6

zepdrix (zepdrix):

?? 0_o Why did your 8 and 6 switch places?

OpenStudy (anonymous):

because i used\[d/dx (u^n)\]

OpenStudy (anonymous):

=nu^n-1

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}(x^n)\quad=\quad nx^{n-1}\]So the 8 comes down, good. Then your 8(exponent) decreases to a 7... yes?

OpenStudy (anonymous):

or i think it should be 48x^7

zepdrix (zepdrix):

yah that looks better.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large\bf\sf h(x)=6x^8f(x)\] \[\Large\bf\sf h'(x)\quad=\quad \color{royalblue}{\left[6x^8\right]'}f(x)+6x^8\color{royalblue}{\left[f(x)\right]'}\]Did you remember about product rule? Here is what the setup should look like. Where the blue parts are the derivatives you need to take.

OpenStudy (anonymous):

\[6x^ 8f(x)+48x^7f(x)\]

OpenStudy (anonymous):

\[48x^7f(x)+6x^8f(x)\]

zepdrix (zepdrix):

Take the derivative of the second f. Your result looks close,\[48x^7f(x)+6x^8f'(x)\]

OpenStudy (anonymous):

\[48^7f(x)+6x^8f\]

zepdrix (zepdrix):

? :\

OpenStudy (anonymous):

\[48x^7f(x)+6x^8f\]

OpenStudy (anonymous):

i don't really understand how to take the derivative of the second f

zepdrix (zepdrix):

Oh boy... The f is NOT a constant coefficient multiplying an x. It's a function of x. The derivative of f(x) is f'(x). You stop there.... We don't KNOW what f(x) is, we can't go any further in defining its derivative.

OpenStudy (anonymous):

so if the f'(x) is constant then it will be \[48x^7f(x)+6x^8\]

zepdrix (zepdrix):

If for some reason they had told us that f'(x) was constant, then we could write it as:\[48x^7f(x)+6x^8\cdot C\]Where C is our constant derivative. They didn't tell us that though :o

OpenStudy (anonymous):

ok so your saying that the f(x) derivative is f'(x), so therefor my answer should be \[48x^7f(x)+6x^8f'(x)\]

zepdrix (zepdrix):

yes.

OpenStudy (anonymous):

ok i put that at first but you told me to look at the problem again:(

OpenStudy (anonymous):

oh no i didn't sorry

zepdrix (zepdrix):

ya yours was close :c maybe was just a typo before.

OpenStudy (anonymous):

ok so now i have to plug it into h'(x) right

zepdrix (zepdrix):

No. You have finished part b).\[\Large\bf\sf h(x)\quad=\quad 6x^8\;f(x)\]\[\Large\bf\sf h'(x)\quad=\quad 48x^7\;f(x)+6x^8\;f'(x)\]

zepdrix (zepdrix):

Unless that's what you meant by plug it in.. :o

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large\bf\sf h(x)=\frac{(1-e^x)^3}{f(x)}\]Oh boy, here's where they start to get tricky :D

OpenStudy (anonymous):

lol....ok i'm about to try

OpenStudy (anonymous):

ok can you tell me if i'm on the right track? \[f(x)3(1-e^x)^2-(1-e^x)^3f'(x)/(f(x))^2\]

zepdrix (zepdrix):

Yesssss nice job, you're very close. Just need to fix this orange part a sec,\[\Large\bf\sf h'(x)\quad=\quad \frac{f(x)\cdot\color{orangered}{3(1-e^x)^2}-(1-e^x)^3\;f'(x)}{f(x)^2}\]

zepdrix (zepdrix):

Need to apply the chain rule, not just the power rule to the outside.

zepdrix (zepdrix):

Multiply by the derivative of the inside.

OpenStudy (anonymous):

\[\frac{ f(x)3(1-e^2x)^2-(1-e^x)^3f'(x) }{f(x)^2 }\]

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}(1-e^x)^3\quad=\quad 3(1-e^x)^2\color{royalblue}{\frac{d}{dx}(1-e^x)}\]Chain rule^ This is what you should be getting for that orange part,\[\Large\bf\sf \color{orangered}{3(1-e^x)^2(-e^x)}\]

OpenStudy (anonymous):

ok so that will be \[\frac{ f(x)3(1-e^x)^2(-e)-(1-e^x)^2f'(x) }{ f(x) ^2}\]

zepdrix (zepdrix):

Don't forget the x exponent on your e! :O And you changed the second (1-e^x) term.. hmm it was correct before. Should get something like this,\[\Large\bf\sf h'(x)\quad=\quad \frac{f(x)\cdot\color{orangered}{3(1-e^x)^2(-e^x)}-(1-e^x)^3\;f'(x)}{f(x)^2}\]

OpenStudy (anonymous):

ok now i factor

zepdrix (zepdrix):

I would ummmmm.. leave it like that. You can maybe simplify it a little bit, but I wouldn't fuss with it too much. Just bring the `negative` and `3` to the front.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hello Zepdrix, can you help me with this last problem i'm stuck on

zepdrix (zepdrix):

\[\Large\bf\sf H(x)=5\sqrt {f(x)^4}\]

zepdrix (zepdrix):

Did I format that correctly? The 4th power is inside of the root like that?

OpenStudy (anonymous):

yes ok this is what I have so far\[1/5(f(x)^4)^-4/5*(f(x)^4)\]

OpenStudy (anonymous):

the-4/5 is an exponent

zepdrix (zepdrix):

Oh is it supposed to be a 5th root? Is that what the 5 is about?

zepdrix (zepdrix):

\[\Large\bf\sf H(x)=\sqrt[5]{f(x)^4}\]

zepdrix (zepdrix):

\[\Large\bf\sf H(x)=f(x)^{4/5}\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\bf\sf H'(x)=\frac{4}{5}f(x)^{-1/5}\cdot f'(x)\]

zepdrix (zepdrix):

I'm a little confused by your coefficient and exponent, they look backwards. Hmm

OpenStudy (anonymous):

ok I started off like this\[h(x)=(f(x)^4)^1/5\]

zepdrix (zepdrix):

Hmm you should simplify it a little further before differentiating. Looks like you're giving yourself more work D: \[\Large\bf\sf \sqrt[5]{a^{4}}\quad=\quad (a^4)^{1/5}\quad=\quad a^{4/5}\]

zepdrix (zepdrix):

Otherwise, from where you left off,\[1/5(f(x)^4)^{-4/5}*\color{orangered}{(f(x)^4)}\]You still need to chain rule this, applying the power rule, then chain rule again to get your f'(x) then a bunch of simplification. It will get pretty messy :\

OpenStudy (anonymous):

\[\frac{ 1 }{ 5}(f(x)^4)^{-4/5}(4f'(x))\]

OpenStudy (anonymous):

then I put \[\frac{ 4f'(x) }{5 \sqrt[5]{f(x)^4} }\]

zepdrix (zepdrix):

\[\frac{ 1 }{ 5}(f(x)^4)^{-4/5}\color{orangered}{(4f(x)^3)f'(x)}\]

zepdrix (zepdrix):

You forgot your power rule.

zepdrix (zepdrix):

This approach is so messyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyy Do it the other way :( Ugh... there is so much room for error doing it this way....

zepdrix (zepdrix):

You're doing an extra chain, and 2 extra simplification steps...

OpenStudy (anonymous):

ok sorry but this is kind of the only way I know

zepdrix (zepdrix):

\[\Large\bf\sf H(x)\quad=\quad \sqrt[5]{f(x)^4}\quad=\quad\left(f(x)^4\right)^{1/5}\quad=\quad \color{orangered}{f(x)^{4/5}}\]That's the only step you're missing. Writing it as a rational expression (fraction for your exponent).\[\Large\bf\sf H(x)\quad=\quad f(x)^{4/5}\]

zepdrix (zepdrix):

Your power rule and chain will work out nicer from there. :o

OpenStudy (anonymous):

ok thankyou I understand now